# Uncertainty of Thoughts: Uncertainty-Aware Planning Enhances Information Seeking in Large Language Models

### v2

@@ 5 Conclusion and Discussion
This paper presents the Uncertainty of Thoughts (UoT) algorithm, significantly improving LLMs in tasks requiring active information seeking through tree-based simulation, uncertainty-based rewards and a reward propagation scheme. On five datasets UoT increases success rate by 38.1% on average, establishing a new benchmark for evaluating LLMs in active information-seeking tasks. We evaluate UoT on simplified scenarios; more realistic scenarios raise challenges like allowing incomplete elimination of possibilities by answers, and others which we leave for future work. We further discuss these limitations and future work in Appendix H.
@@ Appendix E Effect of Simulation Depth
As Figure 5 illustrates, we analyze the impact of simulation steps. Even with one-step reasoning and planning, our method can still have a strong performance, further indicating the effectiveness of our reward design and question selection mechanism. With the increase of the step, the performance can gradually rise. However, due to the constraints of computation resources and OpenAI API budgets, we only explore the simulation to the third step and argue that it can be the practical tradeoff between performance and efficiency.
Figure 5: Results comparison of different depth settings in UoT.
@@ Appendix F Reliability of GPT-4 as the Environment
As the impressive understanding ability of LLMs, previous research has validated the effectiveness of evaluators served by ChatGPT or GPT-4 [7, 17]. Consequently, we also adopt GPT-4 as the environment to provide feedback on our work. Prompts can be found in Appendix L.4. To assess the accuracy and reliability of employing GPT-4 as the environment simulator, we randomly sample 10% interaction records (including the final judgment and intermediate feedback from the environment) from each dataset. As Figure 6 shows, GPT-4 can provide completely accurate judgment and also keep a high level of accurate feedback during the interaction. These experimental results can further support the effectiveness of our method.
@@ Appendix G Reward Function Details and Its Curve
Appendix H Limitation and Future Work
@@ I.2 Scenarios Settings and Datasets
20 Questions game is a classic guessing game where the answerer thinks of an object, person, place, or other, and the questioner, possessing no prior knowledge about the chosen entity, proceeds to pose a series of up to 20 yes-or-no questions to determine what the secret item is. The questions are designed to narrow the possibilities and ultimately guess the secret item within the 20 questions. 20 Questions in BIG-bench: It is the sub-task of BIG-bench and can be found on the GitHub website [^6] https://github.com/google/BIG-bench/tree/main/bigbench/benchmark_tasks/twenty_questions, consist of 29 items. Common Dataset Construction: We came across an official website [^7] https://blog.prepscholar.com/20-questions-game that introduces a 20 Questions game, which mentions that common target categories in this game include animals, places, food, and objects. Therefore, we extracted and manually screened the targets mentioned on this website, resulting in a dataset named "Common" comprising 111 targets, each belonging to one of the four aforementioned categories. Thing Dataset: It is a collection of 1,854 varied object concepts, carefully selected from tangible and easily identifiable nouns in American English by Martin at al.[13], which is publicly available on their official website [^8] https://osf.io/jum2f.
@@ I.3 Dataset selection criteria and process for MedDG
I.4 UoT (Open Set) Setup
@@ I.5 Implementation
Some examples of conversation in different scenarios are provided in Table 7-10.
@@
The specific prompts for the DP(OS) and DP(CS) baselines in each scenario are provided in Table 11-13.
@@
The specific prompts for Planning Prompt baselines in each scenario are provided in Table 14-16. As planning prompt method is close set setting, hence the probability set will also be informed in the prompt as DPCS prompt. We do not repeat it in the tables.
@@
The detailed prompts for our UoT method in each scenario are attached in Table 17-19.
@@
The specific prompt to guide the examiner in all scenarios is provided in Table 20-22.

### v3

@@ 6 Conclusion and Discussion
This paper presents the Uncertainty of Thoughts (UoT) algorithm, significantly improving LLMs in tasks requiring active information seeking through tree-based simulation, uncertainty-based rewards and a reward propagation scheme. On five datasets UoT increases success rate by 38.1% on average, establishing a new benchmark for evaluating LLMs in active information-seeking tasks. We evaluate UoT on simplified scenarios; more realistic scenarios raise challenges like allowing incomplete elimination of possibilities by answers, and others which we leave for future work.
7 Acknowledgment
Pang Wei Koh is supported by the Singapore National Research Foundation and the National AI Group in the Singapore Ministry of Digital Development and Innovation under the AI Visiting Professorship Programme (award number AIVP-2024-001).
@@ Appendix E Effect of Simulation Depth
As the below figure illustrates, we analyze the impact of simulation steps. Even with one-step reasoning and planning, our method can still have a strong performance, further indicating the effectiveness of our reward design and question selection mechanism. With the increase of the step, the performance can gradually rise. However, due to the constraints of computation resources and OpenAI API budgets, we only explore the simulation to the third step and argue that it can be the practical tradeoff between performance and efficiency.
111122223333606060606565656570707070757575758080808085858585909090909595959566.16775.282.686.792.563.564.767.3DepthSuccess Rate (%)20 QuestionsMedical DiagnosisTroubleshooting
@@ Appendix F Reliability of GPT-4 as the Environment
As the impressive understanding ability of LLMs, previous research has validated the effectiveness of evaluators served by ChatGPT or GPT-4 Chiang & Lee (2023); Liu et al. . Consequently, we also adopt GPT-4 as the environment to provide feedback on our work. Prompts can be found in Appendix L.4. To assess the accuracy and reliability of employing GPT-4 as the environment simulator, we randomly sample 10% interaction records (including the final judgment and intermediate feedback from the environment) from each dataset. As Figure 6 shows, GPT-4 can provide completely accurate judgment and also keep a high level of accurate feedback during the interaction. These experimental results can further support the effectiveness of our method.
@@ Appendix G Reward Function Details and Its Curve
Appendix H Experimental Statistical Significance
We conduct three experiments on five datasets using Llama 3 and GPT-4 to compare the performance of Direct Prompting (DP) and UoT methods in a closed-set setting for significance testing. Due to LLM API quota limitations, the number of experiments are restricted. To determine whether the differences in success rates (SR) between the two methods were statistically significant, we performed a t-test. The results are presented below.
GPT-4 Results
Table 7: GPT-4 Comparison of DP and UoT on Success Rates (SR)
Dataset DP (%) UoT (%) t-Statistic p-Value Significance Conclusion
Common 49.0 70.9 -10.8 0.00041 Significant (p $<$ 0.05)
Thing 30.8 36.8 -8.04 0.00129 Significant (p $<$ 0.05)
DX 89.4 97.0 -3.11 0.03581 Significant (p $<$ 0.05)
MedDG 74.9 87.9 -7.33 0.00185 Significant (p $<$ 0.05)
FloDial 42.5 67.8 -19.8 0.00004 Significant (p $<$ 0.05)
Llama 3 Results
Table 8: Llama 3 Comparison of DP and UoT on Success Rates (SR)
Dataset DP (%) UoT (%) t-Statistic p-Value Significance Conclusion
Common 47.7 56.5 -4.39 0.01180 Significant (p $<$ 0.05)
Thing 14.8 24.8 -16.0 0.00009 Significant (p $<$ 0.05)
DX 80.1 90.1 -4.65 0.00966 Significant (p $<$ 0.05)
MedDG 61.3 64.6 -4.15 0.01426 Significant (p $<$ 0.05)
FloDial 29.9 46.4 -10.5 0.00047 Significant (p $<$ 0.05)
The t-test results indicate that UoT significantly outperform DP five datasets (p < 0.05), as evidenced by their higher mean scores.
@@ I.2 Scenarios Settings and Datasets
20 Questions game is a classic guessing game where the answerer thinks of an object, person, place, or other, and the questioner, possessing no prior knowledge about the chosen entity, proceeds to pose a series of up to 20 yes-or-no questions to determine what the secret item is. The questions are designed to narrow the possibilities and ultimately guess the secret item within the 20 questions. 20 Questions in BIG-bench: It is the sub-task of BIG-bench and can be found on the GitHub website [^6] https://github.com/google/BIG-bench/tree/main/bigbench/benchmark_tasks/twenty_questions, consist of 29 items. Common Dataset Construction: We came across an official website [^7] https://blog.prepscholar.com/20-questions-game that introduces a 20 Questions game, which mentions that common target categories in this game include animals, places, food, and objects. Therefore, we extracted and manually screened the targets mentioned on this website, resulting in a dataset named "Common" comprising 111 targets, each belonging to one of the four aforementioned categories. Thing Dataset: It is a collection of 1,854 varied object concepts, carefully selected from tangible and easily identifiable nouns in American English by Martin at al.Hebart et al. (2019), which is publicly available on their official website [^8] https://osf.io/jum2f.
@@ I.3 Dataset selection criteria and process for MedDG
I.4 Data Preprocessing of FloDial
We process the dataset, FloDial, to convert troubleshooting flowcharts into a set of troubleshooting faults. The dialogue is grounded in specific faults, which correspond to the leaf nodes (descriptions and solutions of faults) in the flowcharts. After reviewing all the leaf nodes, we identify 153 faults that had corresponding dialogues. We then use GPT-4 to generate a clear name for each fault based on the descriptions and solutions of faults, and randomly selected one corresponding dialogue history to construct the current dataset.
I.5 UoT (Open Set) Setup
@@ I.6 Implementation
Some examples of conversation in different scenarios are provided in Table 9-12.
@@
The specific prompts for the DP(OS) and DP(CS) baselines in each scenario are provided in Table 13-15.
@@
The specific prompts for Planning Prompt baselines in each scenario are provided in Table 28-27. As planning prompt method is close set setting, hence the probability set will also be informed in the prompt as DPCS prompt. We do not repeat it in the tables.
@@
The detailed prompts for our UoT method in each scenario are attached in Table 19-21.
@@
The specific prompt to guide the examiner in all scenarios is provided in Table 22-24.
@@
Table 25: ToT Prompt for 20 Question game. [C1] is the count of questions asked. The ‘information gained’ marked blue represents the previous interaction history.
Standard Prompt
You are playing the game of 20 questions. I am impersonating the thing, X. You will ask me up to 20 questions that start with ’Is X’ and can only be answered by yes or no, and I will answer each one truthfully based on being X.
[C1] questions have been asked. And now we know:
[information gained]
Design a question about X and can only be answer by YES or NO.
Additional Reminder in Q14 - Q20
Same as prompts in Appendix L.3
Table 26: ToT Prompt for Medical Diagnosis. [C1] is the count of questions asked. The ‘information gained’ marked blue represents the previous interaction history.
Standard Prompt
You are a doctor and your patient reports that: [self-report]
[C1] questions have been asked. And now we know:
[information gained]
Design a question to ask your patient with symptoms about disease and can only be answered by YES or NO.
Additional Reminder in Q14 - Q20
Same as prompts in Appendix L.3
Table 27: ToT prompt for Troubleshooting. [C1] is the count of questions asked. The ‘information gained’ marked blue represents the previous interaction history.
Standard Prompt
You are a technician and your client reports that: [self-report]
[C1] questions have been asked. And now we know:
[information gained]
Design a question to ask your client with specific situation and can only be answered by YES or NO.
Additional Reminder in Q14 - Q20
Same as prompts in Appendix L.3
Table 28: CoT Prompts.
Prologue
Same as prompts in Appendix L.1
Prompt for Generating Question and Explanation
What’s your next question? Let’s think step-by-step and reply me with your explanation.
Your answer should be:
Explanation: [insert step-by-step analysis here]
Question: [next question]
Additional Reminder in Q14 - Q20
Same as prompts in Appendix L.3